1. (a) Find the Cartesian equation of the curve C traced out by a point whose coordinates, in
terms of a parameter θ, are (a cos θ, b sin θ). Obtain the equations of the tangents at θ = θ1
and θ = θ1 + π/2.
Find the coordinates of the points of intersection of the two tangents, and deduce the
Cartesian equation of its locus.
(b) Find the equations of the tangents to the hyperbola x^2 − 9y^2 = 9 that are drawn from the
point (3, 2) and the points of contact. Find the area of the triangle that these tangents
form with their chord of contact
1
Expert's answer
2021-09-15T02:58:37-0400

Given  curve constrained  by  point  acos θ , bsin θ

Now "x = a cos \u03b8 ; y = b sin \u03b8 \u21d2 x a = cos \u03b8 , y b = sin \u03b8 \u21d2 x a ^2 + y b ^2 = 1" This is equation of ellipse

Now at point "\theta"

"(x_1,y_1) =(acos \theta, b sin \theta)"

Hence the equation of the tangent

"\frac{acos \theta_1 x}{a^2}+\frac{bsin \theta_1 y}{b^2}=1\\
\frac{xcos \theta_1 x}{a}+\frac{ysin \theta_1}{b}=1"


Now the equation of tangent at point "\theta ;\theta_1 +\frac{\pi}{2}"

"(a \cos \theta, b \sin \theta) = (a \cos(\theta_1 +\frac{\pi}{2}), b \sin (\theta_1 +\frac{\pi}{2}))\\
= (-a \sin \theta_1, b \cos\theta_1)"

Now equation of tangent

"-\frac{x\sin \theta_1}{a}+\frac{y\cos \theta_1}{b}=1"


The intersection point of both tangents

"x= (1- \frac{u \sin \theta}{b}) \frac{a}{\cos \theta_1}\\
-\frac{\sin \theta_1}{a}(1- \frac{u \sin \theta}{b}) \frac{a}{\cos \theta_1}+\frac{y \cos \theta _1}{b}=1\\
y= b(\sin \theta_1 +\cos \theta_1)\\
Similarly \\
x= a(\cos \theta_1 -\sin\theta_1)\\"

Hence intersection point

"x= a(\cos \theta_1 -\sin\theta_1)\\
y= b(\sin \theta_1 +\cos \theta_1)\\"


Locus of intersection point

"\frac{x^2}{a^2}+\frac{y^2}{b^2}=2"